OpenStudy (anonymous):

Find the limit of the function algebraically. lim_(x->0) ((x^2+3)/x^4)

OpenStudy (anonymous):

$$\frac{x^2+3}{x^4}=\frac{x^2}{x^4}+\frac3{x^4}=\frac1{x^2}+\frac3{x^4}$$you can't simplify this further; as \(x\to0\) both are going to blow up like so: